One of the authors with too much time on his hands weighed each chocolate candy in a bag of 470 plain chocolate candies. Complete parts (a) and (c).
a. One of the chocolate candies weighed 0.776 gram and it was heavier than 24 of the other chocolate candies. What is the percentile of this particular value?
It is the __th percentile. (Round down to the nearest integer as needed).
b. One of the chocolate candies weighed 0.876 gram and it was heavier than 323 of the other chocolate candies. What is the percentile of this particular value?
It is the __th percentile. (Round down to the nearest integer as needed).
c. One of the chocolate candies weighed 0.856 gram and it was heavier than 227 of the other chocolate candies. What is the percentile of this particular value?
It is the __th percentile. (Round down to the nearest integer as needed).


Answer by stanbon(57352)

One of the authors with too much time on his hands weighed each chocolate candy in a bag of 470 plain chocolate candies. Complete parts (a) and (c).
a. One of the chocolate candies weighed 0.776 gram and it was heavier than 24 of the other chocolate candies. What is the percentile of this particular value?
---------
24/470 = 0.051...
It is the 5th percentile. (Round down to the nearest integer as needed).
--------------------------------
b. One of the chocolate candies weighed 0.876 gram and it was heavier than 323 of the other chocolate candies. What is the percentile of this particular value?
---
323/470 = 0.687...
It is the 68th percentile. (Round down to the nearest integer as needed).
--------------------------------
c. One of the chocolate candies weighed 0.856 gram and it was heavier than 227 of the other chocolate candies. What is the percentile of this particular value?
---
227/470 = 0.4829..
It is the 48th percentile. (Round down to the nearest integer as needed).
===========================
Cheers,
Stan H.
==========